In an experiment involving four different groups, each consisting of 8 subjects, the critical F occupies the cell intersected by column and row degrees of freedom, respectively, of

3 and 32.

4 and 32.

4 and 24.

3 and 28.
Answer #1

Number of groups k = 4

Total number of sample n = 4 * 8 = 32

Degrees of freedom = k-1 , n-k = 4-1 , 32-4 = 3 , 28